OpenStudy (helenredk):

determine the answer that established the identity. tan theta/sin theta sec theta =

Nnesha (nnesha):

\[\huge\rm \frac{ \tan \theta }{ \sin \theta \sec \theta }\] like this ?

OpenStudy (helenredk):

yes

Nnesha (nnesha):

alright. rewrite tan and sec in terms of sin,cos

OpenStudy (helenredk):

i hate this stuff. its confusing

Nnesha (nnesha):

what's the inverse of tan ?? what's the definition of cos, sin tan ?

Nnesha (nnesha):

you should bookmark this website for trig ratios and identity http://www.analyzemath.com/trigonometry/trigonometric_formulas.html look at number 5 (Relations Between Trigonometric Functions) )

OpenStudy (helenredk):

thank you

Nnesha (nnesha):

 tan theta = sinx/cosx and secx=1/cosx so \[\large\rm \frac{\frac{\sin \theta}{\cos \theta}}{\sin \theta \cdot \frac{1}{\cos \theta}}\] now you can change division to multiplication multiply tthe top fraction with the reciprocal of the bottom fraction